Ticket Number: HEA-NOC/20060524-7 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20060524 19:12 UTC+1 Problem Start: 20060526 09:00 UTC+1
Ticket Update: 20060526 14:33 UTC+1 Problem End: 20060526 17:00 UTC+1
Site/line: Network Scanning, News
There will be outages of the HEAnet usenet service (news.heanet.ie) and the
security scanning service (keyhole.heanet.ie) to facilitate the physicical
migration of the servers to the blanchardstown PoP.
20060524 19:10 CMC Issuing ticket.
20060524 19:12 CMC Updating ticket to reflect correct distribution.
20060526 14:30 SY Usenet news service (marathon) and security scanning
service (keyhole) services and servers relocated physically to new data
centre. Services restored successfully, ticket open for monitoring.
This ticket can be monitored at http://www.hea.net/tickets/20060524-7